OpenStudy (anonymous):

How long does it take for an investment to triple in value if it is invested at 10 % compounded quarterly?

OpenStudy (foolaroundmath):

Do you know how to calculate the compound interest?

OpenStudy (anonymous):

A = P(1 + r/n)nt

OpenStudy (foolaroundmath):

Yes, that is correct. Now what is A, r, n and t in that? btw, in your question is the rate of interest 10% per annum (or per quarter)?

OpenStudy (anonymous):

I think it might be per annum because it is asking for a number of years as the answer

OpenStudy (foolaroundmath):

Okay, I'll take the 10% to be per annum. In the formula \(\displaystyle A = P\left ( 1 + \frac{r}{n} \right)^{nt}\), \(A\) is the amount after \(t\) years, \(r\) is the rate of interest per annum and \(n\) is the number of times the interest is calculated in a year. \(t\) is the number of years required.

OpenStudy (anonymous):

r=0.10, n=4

OpenStudy (anonymous):

I don't know what A or t would be

OpenStudy (foolaroundmath):

Yes that's good for now.

OpenStudy (foolaroundmath):

Now, you want the total value to be three times of the invested amount. Here the invested amount is \(P\). And the amount = total value = \(3 \times P\) Make sense ?

OpenStudy (anonymous):

Yeah that makes sense

OpenStudy (foolaroundmath):

great. Now, you want to find out how many years it will take to get that Let's plug in the values now: \(A = 3P\), \(r = 0.1\) and \(n = 4\) to get \(\displaystyle 3\cancel{P} = \cancel{P}\left(1 + \frac{0.1}{4}\right)^{4t} \) \(\displaystyle \left(1 + 0.025\right)^{4t} = 3\) or \( 1.025^{4t} = 3\) everything fine till now?

OpenStudy (anonymous):

Yes that looks good

OpenStudy (foolaroundmath):

Okay, now the major problem comes because that \(t\) is a power. So, you can take the \(\log\) on both sides to make it easier! \(\log(1.025)^{4t} = \log(3)\) Now \( \log(a^b) = b \log (a)\), using this we can simplify the equation to \(4t \times \log(1.025) = \log(3)\) Can you solve the rest now? using a calculator of course ;)

OpenStudy (anonymous):

I ended up with 11.12 years

OpenStudy (anonymous):

I got .0429t=.4771 and divided .4771 by .0429

OpenStudy (foolaroundmath):

That seems to be right :)

OpenStudy (anonymous):

Awesome, thank you!

OpenStudy (foolaroundmath):

You're welcome!